OpenStudy (anonymous):

What is 1and 7 over10 written as an improper fraction?

OpenStudy (kuoministers):

Hmm... 1 7/10 means there is one set of 10 and there is 7 left over Which gives you 17/10

OpenStudy (cwrw238):

multiply 1 by 10 and add 7 to get the numerator givivg 17/10

OpenStudy (cwrw238):

an other example 2 2/3 numerator = 2* 3 + 2 = 8 giving 8/3